OpenStudy (alante):

Solve the equation: |2x-10|=26

jagr2713 (jagr2713):

hey ok what do you think we do first @alante

OpenStudy (alante):

@jagr2713 idk what to do im lost

jagr2713 (jagr2713):

ok well first we can get this into 2 answers so thats good. so we break down the problem into these 2 equations 2x−10=26 −(2x−10)=26 can you solve the first one?

OpenStudy (alante):

(-2x+10)=26 @jagr2713 that it

jagr2713 (jagr2713):

no the firs tone 2x-10=26 what do we do first? hint: its with the -10

OpenStudy (alante):

@jagr2713 turn -10 to 10

jagr2713 (jagr2713):

seems like you're having trouble so i will help So when we have an equation like this 2x-10=26 the first thing we do is subtract or add on both sides so we have -10 so it turns to positive and we add except for subctract but if it was positive then we would subtract. So 10+26=36 so we have 2x=36 so now all we have to do is divide and get x=18 so now we have x by it self. Now try the next one −(2x−10)=26 Hint: we have to distribute first. do you know how to distribute?

OpenStudy (alante):

no @jagr2713

jagr2713 (jagr2713):

Distributing is like this Imagine we have an equation 2(2x-10) so all we do is multiply 2 to all the numbers *inside* the parentheses. 2*2x=4x 2*-10=-20 Like that so now we have for our equation is −(2x−10)=26 so just multiply -1 to all the numbers *Inside* the parentheses -1*2x=? -1*-10=?

jagr2713 (jagr2713):

you there?

OpenStudy (alante):

-2 and 10 @jagr2713

jagr2713 (jagr2713):

you forgot the x so -2x and 10

jagr2713 (jagr2713):

so now we have -2x+10=26 can you solve from here?

OpenStudy (alante):

answer -8 @jagr2713

jagr2713 (jagr2713):

good so we have x=18,-8 as the answr \[ \huge{\color{purple}{\textbf{W}} \color{orange}{\cal{E}} \color{green}{\mathbb{L}} \color{blue}{\mathsf{C}} \color{maroon}{\rm{O}} \color{red}{\tt{M}} \color{gold}{\tt{E}} \space \color{orchid}{\mathbf{T}} \color{Navy}{\mathsf{O}} \space \color{OrangeRed}{\boldsymbol{O}} \color{Olive}{\mathbf{P}} \color{Lime}{\textbf{E}} \color{DarkOrchid}{\mathsf{N}} \color{Tan}{\mathtt{S}} \color{magenta}{\mathbb{T}} \color{goldenrod}{\mathsf{U}} \color{ForestGreen}{\textbf{D}} \color{Salmon}{\mathsf{Y}} \ddot \smile }\] if you want to know more about this site, Here is a tutorial about this site : http://openstudy.com/study#/updates/5525c0fce4b030ceb3f357e8

OpenStudy (alante):

thanks @jagr2713